# SRNN: Spatiotemporal Relational Neural Network for Intuitive Physics Understanding

### v1

@@ Abstract
Human prowess in intuitive physics remains unmatched by machines. To bridge this gap, we argue for a fundamental shift towards brain-inspired computational principles. This paper introduces the Spatiotemporal Relational Neural Network (SRNN), a model that establishes a unified neural representation for object attributes, relations, and timeline, with computations governed by a Hebbian “Fire Together, Wire Together” mechanism across dedicated What and How pathways. This unified representation is directly used to generate structured linguistic descriptions of the visual scene, bridging perception and language within a shared neural substrate. Moreover, unlike the prevalent “pretrain-then-finetune” paradigm, SRNN adopts a “predefine-then-finetune” approach. On the CLEVRER benchmark, SRNN achieves competitive performance. Our analysis further reveals a benchmark bias, outlines a path for a more holistic evaluation, and demonstrates SRNN’s white-box utility for precise error diagnosis. Our work confirms the viability of translating biological intelligence into engineered systems for intuitive physics understanding.
@@ 1 Introduction
In contrast to the prevalent "pretrain-then-finetune" paradigm [baldassarre2025featuresdinofoundationvideo; ICLR2025_5e8309c9; wang2025internvideo25empoweringvideomllms]—where a base model is first pre-trained on massive-scale data and subsequently fine-tuned with limited high-quality samples for downstream tasks—our SRNN adopts a "predefine-then-finetune" approach: it architecturally hard-codes human cognitive priors at the outset. This foundational design allows the model’s parameters to be efficiently tuned from only a handful of samples.
@@ 2 Relative Work
Figure 1: (a) Semantic Nature Design. The core of this design is #action which connects eight semantic roles. (b) Spatial Nature Design. For #change_direction, we add #relation_attr neurons (e.g. #right, #left, #back, etc.) to indicate its attributes. For other relations, #relation_attr does not exist. (c) Connections between spatial neurons and semantic neurons. The neuron #wernicke makes the semantic neurons capable of being activated when language is necessary. All neurons are depicted in blue to indicate their inactive state.
@@
Intuitive Physics Reasoning Task. Neuro-symbolic approaches have been widely explored for this task. NS-DR [yi2019clevrer] represents the spatiotemporal relations of each frame in a video as a directed graph. However, a unified spatiotemporal graph is absent, as NS-DR does not have nodes to represent time or static attributes of an object. Moreover, unlike human brain, where relational concepts are stored in neuronal assemblies [rizzolatti2004mirror], NS-DR uses edges to represent these concepts. DCL [chen2021grounding] and VRDP [chen2021grounding] represent the spatiotemporal relations in each frame of a video as a couple of matrix where each element stands for a numerical confidence score. Object attribute (e.g. shape, color) and relational concepts (e.g. collision) are represented as embedding vectors, thereby being able to optimize a goal function through backpropagation. In terms of representation and computation, the above methods are much different from human brain.
For end-to-end approaches, Object-based Attention [ding2020object] designs a transformer-based neural network for spatio-temporal reasoning about videos. All components are updated by self-supervised learning. In [Lerer2016learning], convolutional neural networks are trained to understand the physical intuition presented in block tower videos. Although the architecture of these neural networks have similarities with human brain, the computational mechanisms underlying neuron connection weights are fundamentally distinct from the Hebbian principle.
@@ 3 SRNN
SRNN first boots spatial and semantic nature designs to predefine its neural network architecture. After that, it begins to watch a video. For each frame of the video, it perceives static states of objects. The attributes of shape, texture and color are attended. In the meanwhile, pixel coordinates of object centers are transformed into 3D camera coordinates by depth estimation. We track these camera coordinates to form object trajectories. Then it perceives spatiotemporal relations between each pair of objects given their trajectories within a time slot. After that, neurons are fired and wired, which occurs simultaneously within three key components: the how pathway for object relations, the what pathway for object attributes, and the temporal binding for timeline construction. A semantic network is also activated and generates sentences to describe the video.
@@ 3.3 Perceiving Spatiotemporal Relations
Figure 2: The Fire-and-Wire Mechanism (left) and the Language Generation Module (right) in the How Pathway. Left: Visual perceptions activate a relational neuron #relation along with creation and activation of entity-instance neurons ins_entity_id1 and ins_entity_id2. Then #relation triggers the creation and activation of a action-stamp neuron stamp_action_id which binds ins_entity_id1 and ins_entity_id2. If relational attributes exists for this relation, the corresponding #relation_attr1 is activated by #relation. A concept-instance neuron ins_concept_id1 is pointed by #relation_attr1 and stamp_action_id. Other neurons below the activation threshold remain inactive (shown in blue). Right: The stamp_action_id neuron triggers ins_action_id, which acts as the starting point of the semantic network. Connections are formed between ins_action_id and its lexical neuron _action_name, as well as with #action. Meanwhile, #relation propagates signals to #action and #semantic_roles via predefined neural pathways in Nature Design. Joint signals from ins_action_id and #relation activates #action, which in turn send signals to all the semantic-role neurons. Only #semantic_role1 and #semantic_role2 are activated upon receiving sufficient signals. Finally, stamp_subaction_id1 wires #semantic_role1 and ins_entity_id1, and stamp_subaction_id2 binds #semantic_role2 and ins_entity_id2.
@@ 3.4 Firing and Wiring Neurons
Inspired by the neurobiological concept that “synchronous firing strengthens synaptic connections” [hebb1949organization], we designed a representation mechanism where SRNN processes visual inputs by activating relevant neurons (Fire) and then establishing connections between these activated neurons (Wire). This process encodes both how entities interact or change over time and what attributes of entities are present in the neural network.
@@
Neurons (e.g., #move, #touch) are activated by the observed corresponding relations. All participant instances involved in these relations are represented by entity-instance neurons which simultaneously receive activation. When a relation neuron fires, it triggers the creation of an action-stamp neuron, which connects all the participating entities of the relation. The left network of Fig. 2 details this process.
Figure 3: The Fire-and-Wire Mechanism in the Temporal Binding. The temporal neuron stamp_time_id1 marks the origin of the timeline and points to stamp_time_id2. All action-stamp neurons present during this temporal window are associated with stamp_time_id1. As time progresses, new temporal neurons are generated, activated, and linked sequentially to form an ordered chain.
@@ 3.5 Generating Language Description
The language module generates a sentence description for each relation. It contains two components: a semantic network and a sentence generator. The semantic network fulfills three primary functions: (1) assigning lexical labels to objects; (2) assigning verbal labels to relations; and (3) assigning semantic roles like agent and patient to the participants involved in a relation. In this project, the moving object in a relation serves as the agent. We present this semantic network in detail in Fig. 2 and Fig. 4 with yellow dashed lines. The implementation of the semantic network draws inspiration from the approach outlined in [yang2025propnetwhiteboxhumanlikenetwork].
A key distinction lies in our introduction of subaction-stamp neurons to bind semantic roles and instance entities, thereby explicitly defining the relationship between entities and their corresponding roles.
@@
Figure 4: The Fire-and-Wire Mechanism and the Language Generation Module in the What Pathway. The neuron ins_entity_id initiates the creation and activation of stamp_entity_id. Simultaneously, concept-instance neurons are created and activated to encode the entity’s attributes, which are then bound together by stamp_entity_id. These instance neurons, in turn, activate corresponding lexical neurons (indicated by yellow dashed lines), denoted as _ins_name.
@@ 4 Experiments
We first demonstrate SRNN achieves competitive accuracy on the CLEVRER dataset, proving its effectiveness in capturing essential video information. Next, a cognitive ablation study is conducted to quantify the causal role of specific relations in video reasoning. This study also serves to audit inherent biases in CLEVRER for the development of a more comprehensive benchmark. Finally, a fine-grained error analysis is performed on incorrectly answered cases, leveraging the intrinsic interpretability of our model to categorize failure patterns and identify limitations.
@@
VRDP [chen2021grounding] 95.7 91.4 94.8 84.3 93.4 96.3 91.9
SRNN 81.0 62.9 84.7 58.0 89.2 90.5 80.0
Table 1: Comparison of question-answering accuracy on CLEVRER. The highest accuracy for each question type is highlighted in boldface. Scores are obtained from the original publications of each model. SRNN demonstrates notably competitive accuracy.
@@ 4.1 Performance Comparison
We compare the performance of SRNN on CLEVRER against a comprehensive set of baselines, including: TVQA+ [lei2018tvqa], Memory [fan2019heterogeneous], IEP (V) [johnson2017inferring], TbD-net (V) [mascharka2018transparency], HCRN [le2020hierarchical], MAC (V+) [hudson2018compositional], NS-DR [yi2019clevrer], DCL [chen2021grounding], Object-based Attention [ding2020object], and VRDP [chen2021grounding]. Among these, VRDP [chen2021grounding] represents the state-of-the-art neuro-symbolic model, while Object-based Attention [ding2020object] is the state-of-the-art end-to-end approach.
@@
Implementation Details. Each video is temporally segmented into five time slots, with each slot corresponding to a one-second duration. The video descriptions generated by our SRNN model are utilized as input for a Large Language Model (LLM) to answer the questions. Specifically, for a given video and a question, the textual descriptions from SRNN are concatenated into a prompt along with the question text. To ensure consistent comprehension between humans and the LLM, critical alignments of the question contexts are incorporated into the prompt. Furthermore, specific solving instructions are added to guide the LLM in reasoning through the problem based on the provided logic. All parameters and prompt alignments are tuned on ten videos from the CLEVRER training split, and the final performance is evaluated on the CLEVRER validation split. The results of key parameters are presented in App. B. YOLO training details are appended in App. A. LLM choice and its prompts are explained in App. E .
Results. As shown in Tab. 1, SRNN delivers a strong performance of 90.5% per option and 80.0% per question for explanatory questions. This result places it as the third-best performing model in our comparison, after Object-based Attention (98.5% per opt., 96.0% per ques.) and VRDP (96.3% per opt., 91.9% per ques.). On descriptive questions, SRNN achieves 89.2% accuracy, a result that is on par with the top-performing models, slightly behind VRDP (93.4%). For Counterfactual questions, our method attains 84.7% per option and 58.0% per question, approaching Object-based Attention (91.4% per opt., 75.6% per ques.). These results collectively affirm SRNN’s effectiveness in capturing and encoding the essential relational information presented in a video, which is necessary for complex spatio-temporal and causal reasoning.
@@ 4.2 Cognitive Ablation
In this section, we perform a cognitive ablation study. Unlike conventional ablation that removes structural components of a model (e.g., layers or neurons), we remove the human-understandable physical concepts (e.g., touch, distance change) presented in the videos, to directly interrogate the causal role of specific relations in video reasoning. Moreover, we audit the biases of the CLEVRER dataset to provide a roadmap for designing a more comprehensive benchmark.
By selectively removing relational neurons in the neural network, such as #touch, corresponding physical concepts can be effectively stripped from the video. The resulting textual descriptions are then fed into LLM to answer questions and evaluate accuracy. The configurations of SRNN and LLM remain identical to those used in Sec. 4.1. Tab. 2 presents the changes in answer accuracy of the model when different types of relations are removed for various question categories.
@@ Appendix D SRNN Examples
Figure 7: Neural network representation for video_12214. In this representation, neurons are color-coded according to their type: nature neurons are shown in red, stamp neurons in gray, time neurons in purple, instance neurons in green, and lexical neurons in yellow. Deactivated neurons are omitted for clarity. Arrows on the connections indicate the direction of signal transmission upon neuron activation.
Figure 8: Language generation module for video_12214. Right: The purple boxes on the right mark the language generation process for the How Pathway. A neuron named ins_action_185 is created and activated by the neural network, serving as the start point of the semantic network. A connection is established between ins_action_185 and #action by the neural network after the birth of ins_action_185, making ins_action_185 being able to send a signal to #action. The neural network also establishes a connection between ins_action_185 and a lexical neuron _touch. Next, two semantic role neurons #subject and #object are activated by the joint signals from #touch and #action. Finally, stamp_subaction_190 is created to bind ins_entity_36 and #subject. Similarly, stamp_subaction_191 is created to bind ins_entity_38 and #object. The sentence generator outputs a sentence “ins_entity_36 touches ins_entity_38” based on this semantic network. Left: The left cyan boxes illustrate the language generation process for the What Pathway. The neuron ins_entity_36 triggers the creation and activation of stamp_entity_54. In the meantime, ins_concept_46, ins_concept_49, etc., are created and activated to record the attributes of this entity, and bound by stamp_entity_54. A phrase “yellow metal cube ins_entity_36” is generated based on this semantic network. Similarly, another phrase “rubber gray cylinder ins_entity_38” is also produced. The sentence generator then merges the outputs from the How Pathway and the What Pathway, ultimately yielding the final sentence: “The yellow metal cube ins_entity_36 touches the rubber gray cylinder ins_entity_38.”
Fig. 6 presents an example of our approach for percepting static properties of objects in a single frame from video_12214. Fig. 7 illustrates the neural network representation corresponding to video_12214. Fig. 8 presents the language generation module for video_12214. Figure 9 presents the textual descriptions for each time slot of video_12214.

### v2

@@ Abstract
Human prowess in intuitive physics remains unmatched by machines. To bridge this gap, we argue for a fundamental shift towards brain-inspired computational principles. This paper introduces the Spatiotemporal Relational Neural Network (SRNN), a model that establishes a unified neural representation for object attributes, relations, and timeline, with computations governed by a Hebbian “Fire Together, Wire Together” mechanism across dedicated What and How pathways. This unified representation is directly used to generate structured linguistic descriptions of the visual scene, bridging perception and language within a shared neural substrate. On the CLEVRER benchmark, SRNN achieves competitive performance, thereby confirming its capability to represent essential spatiotemporal relations from the visual stream. Cognitive ablation analysis further reveals a benchmark bias, outlining a path for a more holistic evaluation. Finally, the white-box nature of SRNN enables precise pinpointing of error root causes. Our work provides a proof-of-concept that confirms the viability of translating key principles of biological intelligence into engineered systems for intuitive physics understanding in constrained environments.
@@ 2 Relative Work
Figure 1: (a) Semantic Nature Design. The core of this design is #action which connects eight semantic roles. (b) Spatial Nature Design. The spatial neurons represent spatial relations between objects, in addition to their intrinsic states. For #change_direction, we add #relation_attr neurons (e.g. #right, #left, #back, etc.) to indicate its attributes. For other relations, #relation_attr does not exist. (c) Connections between spatial neurons and semantic neurons. The neuron #wernicke makes the semantic neurons capable of being activated when language is necessary. All neurons are depicted in blue to indicate their inactive state.
@@
Intuitive Physics Reasoning Task. Neuro-symbolic approaches have been widely explored for this task. NS-DR [yi2019clevrer] represents the spatiotemporal relations of each frame in a video as a directed graph. However, a unified spatiotemporal graph is absent, as NS-DR does not have nodes to represent time or static attributes of an object. Moreover, unlike human brain, where relational concepts are stored in neuronal assemblies [rizzolatti2004mirror], NS-DR uses edges to represent these concepts. DCL [chen2021grounding] and VRDP [NEURIPS2021_07845cd9] represent the spatiotemporal relations in each frame of a video as a couple of matrix where each element stands for a numerical confidence score. Object attribute (e.g. shape, color) and relational concepts (e.g. collision) are represented as embedding vectors, thereby being able to optimize a goal function through backpropagation. In terms of representation and computation, the above methods are much different from human brain.
For end-to-end approaches, Object-based Attention [ding2020object] designs a transformer-based neural network for spatio-temporal reasoning about videos. All components are updated by self-supervised learning. In [Lerer2016learning], convolutional neural networks are trained to understand the physical intuition presented in block tower videos. Although the architecture of these neural networks have similarities with human brain, the computational mechanisms underlying neuron connection weights are fundamentally distinct from the Hebbian principle. Backpropagation conveys global error signals, whereas Hebbian rule relies exclusively on local activity.
@@ 3 SRNN
SRNN first boots spatial and semantic nature designs to predefine its neural network architecture. After that, it begins to watch a video. For each frame of the video, it perceives static states of objects. The attributes of shape, texture and color are attended. In the meanwhile, pixel coordinates of object centers are transformed into 3D camera coordinates by depth estimation. We track these camera coordinates to form object trajectories. Then it perceives spatiotemporal relations between each pair of objects given their trajectories within a time slot. After that, neurons are fired and wired, which occurs simultaneously within three key components: a How pathway for object relations, a What pathway for object attributes, and a temporal binding for timeline construction. A semantic network is also activated and generates sentences to describe the video.
@@ 3.3 Perceiving Spatiotemporal Relations
Figure 2: The Fire-and-Wire Mechanism (left) and the Language Generation Module (right) in How Pathway. Left: Visual perceptions activate a relational neuron #relation along with creation and activation of entity-instance neurons ins_entity_id1 and ins_entity_id2. Then #relation triggers the creation and activation of a action-stamp neuron stamp_action_id which binds ins_entity_id1 and ins_entity_id2. If relational attributes exists for this relation, the corresponding #relation_attr1 is activated by #relation. A concept-instance neuron ins_concept_id1 is pointed by #relation_attr1 and stamp_action_id. Other neurons below the activation threshold remain inactive (shown in blue). Right: The stamp_action_id neuron triggers ins_action_id, which acts as the starting point of the semantic network. Connections are formed between ins_action_id and its lexical neuron _action_name, as well as with #action. Meanwhile, #relation propagates signals to #action and #semantic_roles via predefined neural pathways in Nature Design. Joint signals from ins_action_id and #relation activate #action, which in turn sends signals to all the semantic-role neurons. Only #semantic_role1 and #semantic_role2 are activated upon receiving sufficient signals. Finally, stamp_subaction_id1 wires #semantic_role1 and ins_entity_id1, and stamp_subaction_id2 binds #semantic_role2 and ins_entity_id2.
@@ 3.4 Firing and Wiring Neurons
Inspired by the neurobiological concept that “synchronous firing strengthens synaptic connections” [hebb1949organization], we designed a representation mechanism where SRNN processes the perceived static states and spatiotemporal relations by activating relevant neurons (Fire) and then establishing connections between these activated neurons (Wire). This process encodes both how entities interact or change over time and what attributes of entities are present in the neural network.
@@
Neurons (e.g., #move, #touch) in Spatial Nature Design are activated by the observed corresponding relations. All participant instances involved in these relations are represented by entity-instance neurons which simultaneously receive activation. When a relation neuron fires, it triggers the creation of an action-stamp neuron, which connects all the participating entities of the relation. The left network of Fig. 2 details this process.
Figure 3: The Fire-and-Wire Mechanism in Temporal Binding. The temporal neuron stamp_time_id1 marks the origin of the timeline and points to stamp_time_id2. All action-stamp neurons present during this temporal window are associated with stamp_time_id1. As time progresses, new temporal neurons are generated, activated, and linked sequentially to form an ordered chain.
@@ 3.5 Generating Language Description
The language module generates a sentence description for each relation. It contains two components: a semantic network and a sentence generator. The semantic network fulfills three primary functions: (1) assigning lexical labels to objects; (2) assigning verbal labels to relations; and (3) assigning semantic roles like agent and patient to the participants involved in a relation. In this project, the moving object in a relation serves as the agent. We present this semantic network in detail in Fig. 2 and Fig. 4 with yellow dashed lines. The implementation of the semantic network draws inspiration from the approach outlined in [yang2025propnetwhiteboxhumanlikenetwork]. In contrast to their framework, a key distinction lies in our introduction of subaction-stamp neurons to bind semantic roles and instance entities, thereby explicitly defining the relationship between entities and their corresponding roles.
@@
Figure 4: The Fire-and-Wire Mechanism and the Language Generation Module in What Pathway. The neuron ins_entity_id initiates the creation and activation of stamp_entity_id. Simultaneously, concept-instance neurons are created and activated to encode the entity’s attributes, which are then bound together by stamp_entity_id. These instance neurons, in turn, activate corresponding lexical neurons (indicated by yellow dashed lines), denoted as _ins_name.
@@ 4 Experiments
We first demonstrate SRNN achieves competitive accuracy on the CLEVRER dataset, proving its effectiveness in capturing essential video information. Next, a cognitive ablation study is conducted to quantify the causal role of specific relations and timeline in video reasoning. This study also serves to audit inherent biases in CLEVRER for the development of a more comprehensive benchmark. Finally, a fine-grained error analysis is performed on incorrectly answered cases, leveraging the intrinsic interpretability of our model to categorize failure patterns and identify limitations.
@@
VRDP [NEURIPS2021_07845cd9] 95.7 91.4 94.8 84.3 93.4 96.3 91.9
SRNN + LLM 81.0 62.9 84.7 58.0 89.2 90.5 80.0
Table 1: Comparison of question-answering accuracy on CLEVRER. The highest accuracy for each question type is highlighted in boldface. All baseline scores are sourced directly from [NEURIPS2021_07845cd9]. SRNN demonstrates notably competitive accuracy.
@@ 4.1 Performance Comparison
We compare the performance of SRNN on CLEVRER against a comprehensive set of baselines, including: TVQA+ [lei2018tvqa], Memory [fan2019heterogeneous], IEP (V) [johnson2017inferring], TbD-net (V) [mascharka2018transparency], HCRN [le2020hierarchical], MAC (V+) [hudson2018compositional], NS-DR [yi2019clevrer], DCL [chen2021grounding], Object-based Attention [ding2020object], and VRDP [NEURIPS2021_07845cd9]. Among these, VRDP [NEURIPS2021_07845cd9] represents the state-of-the-art neuro-symbolic model, while Object-based Attention [ding2020object] is the state-of-the-art end-to-end approach.
@@
Implementation Details. Each video is temporally segmented into five time slots, with each slot corresponding to a one-second duration. The video descriptions generated by SRNN are utilized as input for a Large Language Model (LLM) to answer the questions. Specifically, for a given video and a question, the textual descriptions from SRNN are concatenated into a prompt along with the question text. To ensure consistent comprehension between humans and LLM, critical alignments of the question contexts are incorporated into the prompt. Furthermore, specific solving instructions are added to guide LLM in reasoning through the problem based on the provided logic. All parameters and prompt alignments are tuned on ten videos from the CLEVRER training split, and the final performance is evaluated on the CLEVRER validation split. The results of key parameters are presented in App. B. YOLO training details are appended in App. A. We select Qwen3 [qwen3] for this task, and more details about the LLM choice and its prompts are provided in App. E.
Results. As shown in Tab. 1, our method delivers a strong performance of 90.5% per option and 80.0% per question for explanatory questions. This result places it as the third-best performing model in our comparison, after Object-based Attention (98.5% per opt., 96.0% per ques.) and VRDP (96.3% per opt., 91.9% per ques.). On descriptive questions, our method achieves 89.2% accuracy, a result that is on par with the top-performing models, slightly behind VRDP (93.4%). For counterfactual questions, our method attains 84.7% per option and 58.0% per question, approaching Object-based Attention (91.4% per opt., 75.6% per ques.). These results collectively affirm SRNN’s effectiveness in capturing and encoding the essential relational information presented in a video, which is necessary for complex spatio-temporal and causal reasoning.
@@ 4.2 Cognitive Ablation
In this section, we perform a cognitive ablation study. Unlike conventional ablation that removes structural components of a model (e.g., layers or neurons), we remove the human-understandable physical concepts (e.g., touch, distance change) and the timeline presented in the videos, to directly interrogate the causal role of specific relations in video reasoning. Moreover, we audit the biases of the CLEVRER dataset to provide a roadmap for designing a more comprehensive benchmark.
By selectively removing relational neurons in the neural network, such as #touch, corresponding physical concepts can be effectively stripped from the video. For timeline ablation, we randomize the linking order of the time-stamp neurons to simulate the temporal disorder of relational events in the video. The resulting textual descriptions are then fed into LLM to answer questions and evaluate accuracy. The configurations of SRNN and LLM remain identical to those used in Sec. 4.1. Tab. 2 presents the changes in answer accuracy of the model when different types of relations are removed for various question categories.
@@
By treating the LLM as a fixed reasoning engine, the drastic performance drop when removing specific relations proves that the LLM’s success is not inherent but is critically enabled by the information within our representation. Therefore, this study cleanly separates the SRNN’s role as the information provider from the LLM’s role as a powerful reasoner.
@@ Appendix D SRNN Examples
Figure 7: Neural network representation for video_12214 presented in an interactive visualization interface. In this representation, neurons are color-coded according to their type: nature neurons are shown in red, stamp neurons in gray, time neurons in purple, instance neurons in green, and lexical neurons in yellow. Deactivated neurons are omitted for clarity. Arrows on the connections indicate the direction of signal transmission upon neuron activation. Neurons can be dragged to visualize their connections with other neurons.
Figure 8: Language generation module for video_12214. Right: The purple boxes on the right mark the language generation process for the How Pathway. A neuron named ins_action_185 is created and activated by the neural network, serving as the start point of the semantic network. A connection is established between ins_action_185 and #action by the neural network after the birth of ins_action_185, making ins_action_185 being able to send a signal to #action. The neural network also establishes a connection between ins_action_185 and a lexical neuron _touch. Next, two semantic role neurons #subject and #object are activated by the joint signals from #touch and #action. Finally, stamp_subaction_190 is created to bind ins_entity_36 and #subject. Similarly, stamp_subaction_191 is created to bind ins_entity_38 and #object. The sentence generator outputs a sentence “ins_entity_36 touch ins_entity_38” based on this semantic network. Left: The left cyan boxes illustrate the language generation process for the What Pathway. The neuron ins_entity_36 triggers the creation and activation of stamp_entity_54. In the meantime, ins_concept_46, ins_concept_49, etc., are created and activated to record the attributes of this entity, and bound by stamp_entity_54. A phrase “yellow metal cube ins_entity_36” is generated based on this semantic network. Similarly, another phrase “rubber gray cylinder ins_entity_38” is also produced. The sentence generator then merges the outputs from the How Pathway and the What Pathway, ultimately yielding the final sentence: “The yellow metal cube ins_entity_36 touch the rubber gray cylinder ins_entity_38.” Note that the language generation module currently does not include handling of subject-verb agreement.
Fig. 6 presents an example of our approach for perceiving static properties of objects in a single frame from video_12214. Fig. 7 illustrates the neural network representation corresponding to video_12214. Fig. 8 presents the language generation module for video_12214. Figure 9 presents the textual descriptions for each time slot of video_12214.